OpenStudy (anonymous):

Check my answer please... Solve (3x+4)^1/2=4 x = 0 x = 4 x = 6 None of the above i got A

ganeshie8 (ganeshie8):

(3x+4)^1/2=4 is same as \(\large \sqrt{3x+4} = 4\)

ganeshie8 (ganeshie8):

square both sides

OpenStudy (sumi29):

You should get x=4 by squaring both the sides and solving for x.

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

ok i found my mistake thanks for pointing it out to me

OpenStudy (sumi29):

You are welcome. :-)